subtract. simplify if possible
6-x/x-2 - 4x-7/2-x=

Answer by Maths68(1474)

=(6-x)/(x-2) - (4x-7)/(2-x)
=(6-x)/(x-2) - (7-4x)/(x-2)[Multiplying by (-1) both numerator and denominator]
=(6-x)/(x-2) - (7-4x)/x-2)
Taking LCM
=[(6-x) - (7-4x)]/(x-2)
=(6-x-7+4x)/(x-2)
=(3x-1)/(x-2)